A swimmer weighing 60kg dived to a depth of 2 meters in a pool that is 10 m deep. Find the potential energy of the swimmer if we take level 0 of the energy calculation a.) 6 m deep from the surface of the water in the pool. b.) at the bottom of the pool c.) If the swimmer dives to a depth of 4m. What is the change in swimmer potential energy with respect to each level 0 of the calculation.


1
Expert's answer
2021-04-16T07:23:54-0400

a.) 6 m deep from the surface of the water in the pool:


"E=mgh=60·9.8(4)=2352\text{ J}.\\
\Delta E=mg\Delta h=60·9.8(2-6)=-2352\text{ J}."

b.) at the bottom of the pool


"E=mgh=60·9.8(8)=4704\text{ J}.\\
\Delta E=mg\Delta h=60·9.8(2-10)=-4704\text{ J}."

c.) If the swimmer dives to a depth of 4m. (What is the level 0 here? Unclear. That is why assume level 0 is the surface of water)


"E=mgh=60·9.8(-4)=-2352\text{ J}.\\
\Delta E=mg\Delta h=60·9.8(0-(-4))=2352\text{ J}."